Daily samples of 150 sand filters are removed from Pentair Corporation's assembly line and inspected for defects. Over the past 22 days, the following information has been gathered. Develop a (95.45% confidence level) p-chart.

Day Number of Defective Filters Day Number of Defective Filters
1 7 12 9
2 15 13 11
3 11 14 20
4 17 15 21
5 12 16 12
6 4 17 6
7 5 18 13
8 10 19 4
9 6 20 6
10 16 21 7
11 7 22

10

a. What is the p-bar?   

b. What is the standard deviation of the p?   

c. What is the UCLp = , and the LCLp =   

d. Is the process in control? If not, which samples are out of control?
Answer #1

Z VALUE FOR 95.4% CONFIDENCE = 2
SAMPLE SIZE = 150
NUMBER OF SAMPLES TAKEN = 22

P-BAR = TOTAL NUMBER OF DEFECTS / (SAMPLE SIZE * NUMBER OF SAMPLES = 229 / (150 * 22) = 0.069
STDEV = SQRT((PBAR * (1 - PBAR)) / SAMPLE SIZE = SQRT((0.069 * (1 - 0.069)) / 150 = 0.021

UCL = PBAR + (Z * STDEV) = 0.069 + (2 * 0.021) = 0.111
LCL = PBAR - (Z * STDEV) = 0.069 - (2 * 0.021) = 0.027

2. N0, THE PROCESS IS NOT IN CONTROL SINCE SAMPLE 4, 6, 14, 15, 19 ARE OUTSIDE THE CONTROL LIMIT.

FRACTION DEFECTIVE - UCL - LCL - PBAR 0.15 0.1333 0.1133 0.1067 - 0.12 0.0867 0.08 0.0733 0.0667 0.0667 0.06 0.0467 0.0467 0.

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.